# Positive polytopes with few facets in the Grassmannian

### v1

@@ 5 Examples
The quadric surface $Q_{p_{12},p_{34}}=\{qq=p_{12}=p_{34}=0\}$ contains six residual points, and the boundary curve is defined by the four hyperplanes $h_{1}\cdot h_{2}\cdot p_{23}\cdot p_{14}=0$. So the quadratic forms in $\{p_{13},p_{14},p_{23},p_{24}\}$ that vanish in these six residual points intersect the surface $Q_{p_{12},p_{34}}$ in a net of curves that are adjoint to the face in this surface. Furthermore the restriction map is an isomorphism between the net of adjoints to $S$ and this net of curves on the quadric surface. On the two boundary conics defined by $h_{1}$ and $h_{2}$ there are two pairs of boundary vertices. On both $\{h_{1}=0\}$ and $\{h_{2}=0\}$ they are the intersection points of these hyperplanes with the quadrangle $\{p_{12}=p_{34}=p_{14}p_{23}=p_{13}p_{24}=0\}$. As in [10, Example 2.13] and Lemma 5.3 there is a unique (up to sign) adjoint such that the iterated residue of the form $\Omega_{S}$ along the boundary conic both in $\{h_{1}=0\}$ and in $\{h_{2}=0\}$ has residues ±$\pm 1$ at the pairs of boundary vertices. In particular
@@
Now, we have fixed $\Omega_{S}$ to have iterated residues ±$\pm 1$ at the eight boundary points on the two conics $C_{1}=\{p_{12}=p_{34}=h_{1}=0\},C_{2}=\{p_{12}=p_{34}=h_{2}=0\}$ in $\mathrm{Gr}(2,4)$. Each of these eight boundary points is an endpoint of three boundary segments in the lines. One of these segments has one of the endpoints on $C_{1}$ and the other on $C_{2}$. So two segments at each of the eight boundary points have an endpoint among the eight points in $\{h_{1}=h_{2}=0\}$ outside $\{p_{12}=p_{34}=0\}$. These eight points are contained in the eight planes $\{p_{12}p_{34}=p_{14}p_{23}=p_{13}p_{24}=0\}$, one point per plane. These planes each contain a boundary triangle with three boundary vertices. The adjoints of $S$, including the special one $a_{S}$ all vanish on two lines in each of these planes. These lines are also in the algebraic boundary of $S$. The iterated residue of $\Omega_{S}$ along the face $F_{\Pi}$ in such a plane $\Pi$ therefore reduces to a form $\Omega_{F_{\Pi}}$ with no zeros on the plane, and simple poles along the triangle boundary of $F_{\Pi}$. Since $\Omega_{S}$ has iterated residues ±$\pm 1$ at two boundary points of $F_{\Pi}$, so does $\Omega_{F_{\Pi}}$ which therefore has residue ±$\pm 1$ also in the third boundary point. So the iterated residue of $\Omega_{S}$ is ±$\pm 1$ at every boundary point of $S$, and we may conclude, by Lemma 5.1 that $(\mathrm{Gr}(2,4),S)$ is a positive geometry with canonical form $\Omega_{S}$.
Finally, we describe a new phenomenon in this example. In the quadric surface $Q_{h_{1},h_{2}}=\{qq=h_{1}=h_{2}=0\}$ there are two disjoint two-dimensional faces $F_{1}$ and $F_{2}$. The face $F_{1}$ has two boundary conic curves with boundary vertices $\{qq=h_{1}=h_{2}=p_{12}=p_{14}=0\}$. The face $F_{2}$ is bounded by the four conics $\{qq=h_{1}=h_{2}=p_{12}\cdot p_{23}\cdot p_{34}\cdot p_{14}=0\}.$ The face $F_{1}$ has no residual arrangement, while $F_{2}$ has a residual arrangement which is the four points on $Q_{h_{1},h_{2}}\cap\mathcal{R}(S)$ and the two boundary points on $F_{1}$ that are not on $\mathcal{R}(S)$.
The intersection $Q_{h_{1},h_{2}}\cap\mathcal{R}(S)$ is therefore four points. There is a $5$-dimensional space of quadratic forms modulo $Q_{h_{1},h_{2}}$ that contain these four points. Among them there is the $3$-dimensional space forms that are restrictions of adjoints to $S$ and $3$-dimensional space of adjoints to $F_{2}$ that vanish on all six residual points in $Q_{h_{1},h_{2}}$ to the face $F_{2}$.
@@
The iterated residue of $\Omega_{S}$ along $F$ has the form
@@
where $a_{S,F}$ is the restriction of $a_{S}$ to $Q_{h_{1},h_{2}}$. Each of the conics in the boundary of $F$ contains two boundary segments and four boundary points. There are altogether eight boundary points, one in each of the boundary planes of $S$. Notice that we already saw that the iterated residue of $\Omega_{S}$ at each of these points is ±$\pm 1$.
Now, consider the iterated residue of $\Omega_{S}$ on the two connected components $F_{1}$ and $F_{2}$ of $F$. The boundary of $F_{1}$ is the union of two conics $\{qq=h_{1}=h_{2}=p_{12}\cdot p_{14}=0\}$ and two boundary points $\{h_{1}=h_{2}=p_{12}=p_{14}=p_{13}\cdot p_{24}=0\}$. The iterated residue of $\Omega_{S}$ on $F_{1}$ is a form $\Omega_{F_{1}}$ with poles along the two conics, no zeros, and iterated residues ±$\pm 1$ at the two boundary points. In particular $\Omega_{F_{1}}$ is of the form
@@
where $a_{S,F_{2}}$ is an adjoint of $F_{2}$ on $Q_{h_{1},h_{2}}$. It vanishes at the four points of intersection $Q_{h_{1},h_{2}}\cap\mathcal{R}(S)$ between the residual arrangement of $S$ and $Q_{h_{1},h_{2}}$, and at the boundary vertices of $F_{1}$. The iterated residues of $\Omega_{F_{2}}$ at the boundary vertices of $F_{2}$ are ±$\pm 1$, while the residues of this form at the boundary vertices of $F_{1}$ vanish.
@@
and $\lambda=\pm 1$ since the residues are ±$\pm 1$ at the boundary vertices. Similarly, the residue $\Omega_{F_{2}}$ of $\Omega_{S}$ is
@@
becomes the iterated residue of $\Omega_{S}$ along $F$. It has simple poles along the boundary segments and iterated residues ±$\pm 1$ at the boundary points.
@@ 6 Open questions and future directions
We compute residual arrangements of positive polytopes by using cylindrical algebraic decomposition in Mathematica. This is a complicated procedure which prevents us from doing the computation for all $3360$ combinatorial types of positive hexahedra. Is there more efficient software (or a more efficient algorithm) to do this? If not, can one develop such software?

### v2

@@ 5 Examples
The quadric surface $Q_{p_{12},p_{34}}=\{qq=p_{12}=p_{34}=0\}$ contains six residual points, and the boundary curve is defined by the four hyperplanes $h_{1}\cdot h_{2}\cdot p_{23}\cdot p_{14}=0$. So the quadratic forms in $\{p_{13},p_{14},p_{23},p_{24}\}$ that vanish in these six residual points intersect the surface $Q_{p_{12},p_{34}}$ in a net of curves that are adjoint to the face in this surface. Furthermore the restriction map is an isomorphism between the net of adjoints to $S$ and this net of curves on the quadric surface. On the two boundary conics defined by $h_{1}$ and $h_{2}$ there are two pairs of boundary vertices. On both $\{h_{1}=0\}$ and $\{h_{2}=0\}$ they are the intersection points of these hyperplanes with the quadrangle $\{p_{12}=p_{34}=p_{14}p_{23}=p_{13}p_{24}=0\}$. As in [10, Example 2.13] and Lemma 5.3 there is a unique (up to sign) adjoint such that the iterated residue of the form $\Omega_{S}$ along the boundary conic both in $\{h_{1}=0\}$ and in $\{h_{2}=0\}$ has residues $\pm 1$ at the pairs of boundary vertices. In particular
@@
Now, we have fixed $\Omega_{S}$ to have iterated residues $\pm 1$ at the eight boundary points on the two conics $C_{1}=\{p_{12}=p_{34}=h_{1}=0\},C_{2}=\{p_{12}=p_{34}=h_{2}=0\}$ in $\mathrm{Gr}(2,4)$. Each of these eight boundary points is an endpoint of three boundary segments in the lines. One of these segments has one of the endpoints on $C_{1}$ and the other on $C_{2}$. So two segments at each of the eight boundary points have an endpoint among the eight points in $\{h_{1}=h_{2}=0\}$ outside $\{p_{12}=p_{34}=0\}$. These eight points are contained in the eight planes $\{p_{12}p_{34}=p_{14}p_{23}=p_{13}p_{24}=0\}$, one point per plane. These planes each contain a boundary triangle with three boundary vertices. The adjoints of $S$, including the special one $a_{S}$ all vanish on two lines in each of these planes. These lines are also in the algebraic boundary of $S$. The iterated residue of $\Omega_{S}$ along the face $F_{\Pi}$ in such a plane $\Pi$ therefore reduces to a form $\Omega_{F_{\Pi}}$ with no zeros on the plane, and simple poles along the triangle boundary of $F_{\Pi}$. Since $\Omega_{S}$ has iterated residues $\pm 1$ at two boundary points of $F_{\Pi}$, so does $\Omega_{F_{\Pi}}$ which therefore has residue $\pm 1$ also in the third boundary point. So the iterated residue of $\Omega_{S}$ is $\pm 1$ at every boundary point of $S$, and we may conclude, by Lemma 5.1 that $(\mathrm{Gr}(2,4),S)$ is a positive geometry with canonical form $\Omega_{S}$.
Finally, we describe a new phenomenon in this example. In the quadric surface $Q_{h_{1},h_{2}}=\{qq=h_{1}=h_{2}=0\}$ there are two disjoint two-dimensional faces $F_{1}$ and $F_{2}$. The face $F_{1}$ has two boundary conic curves with boundary vertices $\{qq=h_{1}=h_{2}=p_{12}=p_{14}=0\}$. The face $F_{2}$ is bounded by the four conics $\{qq=h_{1}=h_{2}=p_{12}\cdot p_{23}\cdot p_{34}\cdot p_{14}=0\}.$ The face $F_{1}$ has no residual arrangement, while $F_{2}$ has a residual arrangement which is the four points in $Q_{h_{1},h_{2}}\cap\mathcal{R}(S)$ and the two boundary points in $F_{1}$ that are not on $\mathcal{R}(S)$.
The intersection $Q_{h_{1},h_{2}}\cap\mathcal{R}(S)$ is therefore four points. There is a five-dimensional space of quadratic forms modulo $Q_{h_{1},h_{2}}$ that contain these four points. Among them there is the three-dimensional space forms that are restrictions of adjoints to $S$ and three-dimensional space of adjoints to $F_{2}$ that vanish on all six residual points in $Q_{h_{1},h_{2}}$ to the face $F_{2}$.
@@
The iterated residue of $\Omega_{S}$ along $F:=F_{1}\cup F_{2}$ has the form
@@
where $a_{S,F}$ is the restriction of $a_{S}$ to $Q_{h_{1},h_{2}}$. Each of the conics in the boundary of $F$ contains two boundary segments and four boundary points. There are altogether eight boundary points, one in each of the boundary planes of $S$. Notice that we already saw that the iterated residue of $\Omega_{S}$ at each of these points is $\pm 1$.
Now, consider the iterated residue of $\Omega_{S}$ on the two connected components $F_{1}$ and $F_{2}$ of $F$. The boundary of $F_{1}$ is the union of two conics $\{qq=h_{1}=h_{2}=p_{12}\cdot p_{14}=0\}$ and two boundary points $\{h_{1}=h_{2}=p_{12}=p_{14}=p_{13}\cdot p_{24}=0\}$. The iterated residue of $\Omega_{S}$ on $F_{1}$ is a form $\Omega_{F_{1}}$ with poles along the two conics, no zeros, and iterated residues $\pm 1$ at the two boundary points. In particular $\Omega_{F_{1}}$ is of the form
@@
where $a_{S,F_{2}}$ is an adjoint of $F_{2}$ on $Q_{h_{1},h_{2}}$. It vanishes at the four points of intersection $Q_{h_{1},h_{2}}\cap\mathcal{R}(S)$ between the residual arrangement of $S$ and $Q_{h_{1},h_{2}}$, and at the boundary vertices of $F_{1}$. The iterated residues of $\Omega_{F_{2}}$ at the boundary vertices of $F_{2}$ are $\pm 1$, while the residues of this form at the boundary vertices of $F_{1}$ vanish.
@@
and $\lambda=\pm 1$ since the residues are $\pm 1$ at the boundary vertices. Similarly, the residue $\Omega_{F_{2}}$ of $\Omega_{S}$ is
@@
becomes the iterated residue of $\Omega_{S}$ along $F$. It has simple poles along the boundary segments and iterated residues $\pm 1$ at the boundary points.
@@ 6 Open questions and future directions
We compute residual arrangements of positive polytopes by using cylindrical algebraic decomposition in Mathematica. This is a complicated procedure which prevents us from doing the computation for all $1680$ combinatorial types of positive hexahedra. Is there more efficient software (or a more efficient algorithm) to do this? If not, can one develop such software?
Conflict of interest statement
The authors report there are no competing interests to declare.